# Localized Graph-Based Neural Dynamics Models for Terrain Manipulation

### v1

@@ Abstract
Predictive models can be particularly helpful for robots to effectively manipulate terrains in construction sites and extraterrestrial surfaces. However, terrain state representations become extremely high-dimensional especially to capture fine-resolution details and when depth is unknown or unbounded. This paper introduces a learning-based approach for terrain dynamics modeling and manipulation, leveraging the Graph-based Neural Dynamics (GBND) framework to represent terrain deformation as motion of a graph of particles. Based on the principle that the moving portion of a terrain is usually localized, our approach builds a large terrain graph (potentially millions of particles) but only identifies a very small active subgraph (hundreds of particles) for predicting the outcomes of robot-terrain interaction. To minimize the size of the active subgraph we introduce a learning-based approach that identifies a small region of interest (RoI) based on the robot’s control inputs and the current scene. We also introduce a novel domain boundary feature encoding that allows GBNDs to perform accurate dynamics prediction in the RoI interior while avoiding particle penetration through RoI boundaries. Our proposed method is both orders of magnitude faster than naïve GBND and it achieves better overall prediction accuracy. We further evaluated our framework on excavation and shaping tasks on terrain with different granularity. The project page is available at chaoqi-liu.com/scoopbot.
@@ I INTRODUCTION
In this work, we propose a framework for modeling deformable terrains at fine resolutions over large spatial scales, such as a construction site in which a mobile excavator uses 3D mapping to generate a terrain representation. In particular, our method extends the GBND method to handle spatially large-scale and even unbounded terrains which may be represented by graphs of millions of particles that cannot fit in GPU memory. To predict the deformation of the terrain under a candidate interaction in a computationally efficient manner suitable for planning, our method restricts the attention of dynamics prediction to a dynamically-selected local neighborhood. We introduce a learning-based 3D region-of-interest (RoI) proposer that identifies a small subset of particles from the large graph that are likely to move. Assuming particles that are outside the RoI stay static during robot-terrain interaction improves speed by orders of magnitude in addition to higher dynamics prediction accuracy. Our method simultaneously learns the GBND dynamics and the RoI from simulation data, and both are fine-tuned on real data. Moreover, we introduce a novel boundary feature representation to address the problem that naïvely applying GBND to the identified subgraph leads to erroneous predictions. In our experiments, we verify both the efficiency and effectiveness of the proposed framework for manipulating terrains of different materials.
II BACKGROUND AND RELATED WORK
@@
III METHOD
III-A Overview
@@
Our goal is to learn a computationally-efficient model to predict the shape of a terrain after robot-terrain interaction, such that the model can be used in terrain manipulation planning. Specifically, letting $x_{t}$ denote the state of the terrain, we model $\hat{x}_{t+1}=f(x_{t},u_{t})$ where $u_{t}$ is the robot control and $t$ is the time index. We learn fine-grained prediction where the time step is 0.1 sec and $u_{t}$ indicates a small robot end effector displacement. The dynamics model is trained on data recorded during three types of actions (scooping, pushing, and dumping) and we use a unified dynamics model for all actions. Pretraining is performed from simulation data and then the model is fine-tuned on real data.
We choose GBND as our terrain model, which is a particle-based scene representation that provides flexibility in inter-particle interactions, an inductive bias toward locality in interactions, mass conservation, and preserves fine details of deforming solids and fluids [9, 11, 10].
@@
The core challenge addressed by our method is that the terrain is a volume of material with a very high-dimensional state. Our approach assumes that only a small region of interest (RoI) $R_{t}(u_{t})\subset\mathbb{R}^{3}$ of the volume is affected by the interaction. In other words, letting $p_{t}^{(i)}$ denote the $i^{\text{th}}$ particle in the terrain at time $t$, its future state $\hat{p}_{t+1}^{(i)}=p_{t}^{(i)}$ for any point $p_{t}^{(i)}\notin R_{t}(u_{t})$. Our method simultaneously learns $f$ and $R_{t}$. The RoI proposer is formulated as a super-level set of a learned function $g(p,u_{t})$ such that $g(p,u_{t})\geq 0$ if and only if $p\in R_{t}(u_{t})$. Using the RoI, we only predict the motion of a subgraph of the particle graph. Then, given a terrain manipulation goal, a planner is used to find a near-optimal action sequence. The proposed framework is shown in Fig. 2.
@@
Our key innovations within this framework include: 1) learning $f$ and $g$ to minimize GBND memory usage and computation time while retaining high accuracy; and 2) introducing novel node features that enable learning GBND over arbitrary volumetric domains, ensuring that accuracy is preserved at the boundary of $R_{t}$, and that translation invariance is maintained.
III-B Background on Graph Based Neural Dynamics
@@
This section presents background on the GBND formulation, and for more details we refer readers to Refs. [10, 11, 31]. A GBND describes the environment (robot and terrain) as a set of interacting particles, each of which implicitly represents a small volume of matter, and interactions between particles are learned. The dynamics model is a GNN that represents the environment state at time $t$ as a directed graph (digraph) $x_{t}\equiv G_{t}(V_{t},E_{t})$, where $V_{t}\triangleq\{v_{t}^{(i)}\}_{i=1}^{|V_{t}|}$ is the set of nodes, and $E_{t}\triangleq\{e_{t}^{(k)}\}_{k=1}^{|E_{t}|}$ is the set of edges, with each edge $e_{t}^{(k)}$ denoting a potential direct interaction from $v_{t}^{(i)}$ to $v_{t}^{(j)}$. Indirect interactions between particles are modeled in this framework using a message passing algorithm.
@@
In graph construction, each particle $p_{t}^{(i)}$ is instantiated into a node with feature vector $v_{t}^{(i)}$, and nearby particles are instantiated into edges with the feature vector $e_{t}^{(k)}$. The dynamics are defined as follows. Each node and edge are first encoded into a latent representation,
$z_{0}[v_{t}^{(i)}]=f_{V}^{\text{enc}}(v_{t}^{(i)}),\quad z_{0}[e_{t}^{(k)}]=f_{E}^{\text{enc}}(e_{t}^{(k)}).$ (1)
@@
III-C GBND for Large-Scale Terrain Modeling
We now describe how we use GBND to model terrains with high accuracy and low computational cost. A large-scale graph is first constructed through perception input (Fig. 3). For an observed terrain surface map, we instantiate terrain particles $P_{t}=\{p_{t}^{(i)}\}_{i=1}^{N_{p}}$, through the volume of the terrain down to a sufficiently large depth. We also instantiate tool particles sampled on the robot’s end effector $T_{t}=\{p_{t}^{(i)}\}_{i=N_{p}+1}^{N_{p}+N_{t}}$. For each particle, we connect it to the $k$ closest particles that meet a distance threshold through a nearest-neighbor algorithm.
Our baseline node feature implementation is defined as
where $\mathsf{v}_{t-H:t-1}$ is particle $p^{(i)}$’s velocity over the last $H$ time steps, $u_{t}$ is the impulse on the particle directly caused by robot action, and $c$ is the particle class (terrain or tool) embedding. The edge features for an edge $e_{t}^{(k)}$ connecting particles $i$ and $j$ are
We will show in the next section how including boundary features with nodes $v_{t}^{(i)}$ leads to higher accuracy and generalization when predicting over an RoI.
Because the computation time and memory usage of GBND are closely tied to the number of particles and edges in the graph, limiting the graph’s size helps ensure efficiency. So, at each time step $t$, our RoI proposal method defines the volume $R_{t}(u_{t})\subset\mathbb{R}^{3}$ via the implicit function $g(p,u_{t})$. We describe details of our implementation in Section III-E. The terrain particles $p_{t}^{(i)}\in R_{t}(u_{t})$ are considered to be dynamic, so we extract the subgraph corresponding to these particles and then connect them to tool particles. Only the terrain and tool particles in this subgraph undergo node/edge feature instantiation and GBND propagation.
III-D Encoding RoI Boundaries as Node Features
The standard GBND machinery is best suited for tasks with similar training / test distributions, but our setting introduces the complication of unboundedness and varying regions of interest. As an example, past works using GBND for fluid and pile manipulation have used a fixed domain of particles, such that the network could implicitly learn the boundary conditions of a floor and walls [13, 12, 32, 18]. In our case, “boundary conditions” are not strictly-enforced boundaries, but are rather just more material that is unaffected by the current interactions due to energy dissipation.
In effect, particles adjacent to the RoI boundary should “feel” resistance as though the boundary were fixed. With baseline node encoding, these particles will simply pass through the boundary as if there was nothing there. We also experimented with adding “wall particles”, a boundary set of nodes, directly connected to nodes in the RoI, whose latent features are frozen. Not only does this increase memory usage due to massive wall particles, but in our experiments, inner particles were able to penetrate gaps in the sampled wall particles which introduces unnatural artifacts.
Our approach incorporates an estimated geometric normal of each particle within RoI. Normal estimation is performed using standard techniques within the point cloud of RoI particles [33]. In the interior, the normal information is not likely to help much, but normals are likely to be strongly aligned near the boundary. This allows our network to encode the effect that nodes near the RoI boundary should “feel” resistance in the direction of the normal.
This encoding also has the favorable effect of translational invariance, which is that for the overall dynamics model $f$,
Figure 4: Comparing different methods for encoding RoI boundaries in node features. The baseline (no encoding) does not perform well, whereas encoding absolute height ($z$ features) leads to implicitly learning a “floor effect.” This becomes detrimental when the validation set has different heights from the training set. In contrast, our proposed normal features perform well in all cases.
Experiments in Fig. 4 present validation losses of learned GBND models for different sets of node features. As baselines we compare the standard GBND encoding as well as augmenting with particle height ($z$ features).In the in distribution case, the training and validation sets used terrains whose particles were sampled in a fixed $z$ range of [0 cm, 5 cm]. In the out of distribution case, the validation particle sets were randomly shifted in the $z$ axis by randomly sampled offsets from -100 to 100 m. Encoding the $z$ value significantly improves loss in the in-distribution case, demonstrating that the network is implicitly learning a “floor effect”. This implicit relationship is detrimental in the out-of-distribution case. In contrast, our proposed normal encoding leads to better generalization than the baseline both in and out of distribution.
III-E Convolutional Neural Net RoI Proposer
Finally, we describe how we encode and learn our RoI implicit function $g(p,u_{t})$. We first extract the planar tool position and yaw from $u_{t}$ and define a 2D reference frame $T_{\text{tool}}$. We then predict a tool-centered 2D height function $\tilde{g}(x,y)$ such that $g(p,u_{t})=p_{z}-\tilde{g}(T_{\text{tool}}^{-1}(p_{x},p_{y}))$. In other words, points whose $z$ exceeds $\tilde{g}(x,y)$ (in the tool-centered frame) are included in the RoI.
Our approach uses a terrain heightmap and instantiates $\tilde{g}$ as a Convolutional Neural Net (CNN) architecture. A heightmap is used because it is compatible with elevation mapping pipelines and avoids computations over the large number of volumetric particles. The heightmap with grid resolution 2.5 cm is centered about the tool and cropped to a wide enough range (set to 0.4 m $\times$ 0.4 m in our current implementation).
At each time step, particles in the bounding box are squeezed into a 32$\times$32 heightmap, with each pixel is concatenated with a 16D latent representation of the current control input. Our CNN is a U-net-like [34] architecture consisting of three levels of pooling, nonlinear, and upsampling steps with skip layers. The output CNN is a 16$\times$16 depth map representation of $\tilde{g}$.
Algorithm 1 Dynamics with Scale Reduction
procedure Propose($\{p_{t-H:t}^{(i)}\}_{i=1}^{N},u_{t}$)
@@ III-E Convolutional Neural Net RoI Proposer
    Predict $\tilde{g}$ based on $M_{t-H:t},u_{t}$triangleright $R_{t}$ in Sec. III-E
@@
             $ndices ← i$
@@
We train the RoI predictor with a simulation dataset. To identify particles that are moving, we first apply a velocity threshold of 10 cm/s. Noise in the simulator remains a challenge, and we employ connected component filtering to focus on the main bulk of the terrain movement near the scoop tip. The CNN is trained given the particle heightmap as input and the output is a heightmap representation of $g$. We define the target heightmap as a function that separates the moving and non-moving particles. This is defined by first taking the upper heightmap of non-moving particles, then creating a bowl-shaped function by inflating their value proportionally to the distance to the center, which we found empirically encourages training convergence. Then, we take the min of this map with the bottom-up heightmap of the moving particles. For grid cells in the particle heightmap that are non-empty, we apply an $L_{1}$ loss between the CNN output and the target heightmap. For empty grid cells, we set a hinge loss at a target value of 0.5 m, and do not penalize exceeding this value. The pseudocode of GBND coupled with scale reduction is presented in Algo. 1.
III-F Planning Terrain Manipulation Trajectories
Our planner chooses amongst parameterized scooping, pushing, and dumping trajectories to greedily minimize the difference of the predicted terrain shape to a target heightmap. After every trajectory is executed the large-scale terrain shape is updated and the planner is called again. For scooping and pushing, we inherit the penetrate-drag-scoop (PDS) trajectory parametrization widely used in excavation community [1, 35, 7] (Fig. 5), where the scoop maintains 0 roll angle throughout the entire trajectory. This defines a 9D motion parameter space. For dumping, we use a 6D parameter space for the dumping location, yaw angle, pre-dump and post-dump pitch angles. We also maintain the number of in-scoop particles, which is 0 after dump and push. Scooping and pushing are available if the scoop contains no material.
Algorithm 2 Planning with Learned Dynamics
@@ III-F Planning Terrain Manipulation Trajectories
The optimization is based on the sampling-based planner MPPI [36], which uses the scores of prior trajectories to update a sampling distribution over the parameter space $\Omega$. In each iteration, a batch of motion parameters is sampled $\{\omega_{i}\}_{i=1}^{b}$, and these are then converted into corresponding control sequences, $\{u_{i,0:T_{i}}\}_{i=1}^{b}$. For all sequences, the dynamics model is rolled out (in parallel), and scores are evaluated. The planner maintains the best sequence and updates the distribution to focus on areas of the space in which the score is low. Pseudocode is listed in (Alg. 2).

### v2

@@ Abstract
Predictive models can be particularly helpful for robots to effectively manipulate terrains in construction sites and extraterrestrial surfaces. However, terrain state representations become extremely high-dimensional especially to capture fine-resolution details and when depth is unknown or unbounded. This paper introduces L-GBND, a learning-based approach for terrain dynamics modeling and manipulation, leveraging the Graph-based Neural Dynamics (GBND) framework to represent terrain deformation as motion of a graph of particles. Based on the principle that the moving portion of a terrain is usually localized, our approach builds a large terrain graph (potentially millions of particles) but only identifies a very small active subgraph (hundreds of particles) for predicting the outcomes of robot-terrain interaction. To minimize the size of the active subgraph we introduce a learning-based approach that identifies a small region of interest (RoI) based on the robot’s control inputs and the current scene. We also introduce a novel domain boundary feature encoding that allows GBND to perform accurate dynamics prediction in the RoI interior while avoiding particle penetration through RoI boundaries. Our proposed method is both orders of magnitude faster than naïve GBND and it achieves better overall prediction accuracy. We further evaluated our framework on excavation and shaping tasks on terrain with different granularity.
@@ I Introduction
In this work, we propose a framework, L-GBND, for modeling deformable terrains at fine resolutions over large spatial scales, such as a construction site in which a mobile excavator uses 3D mapping to generate a terrain representation. In particular, our method extends the GBND to handle spatially large-scale and even unbounded terrains which may be represented by graphs of millions of particles that cannot fit in GPU memory. To predict the deformation of the terrain under a candidate interaction in a computationally efficient manner suitable for planning, our method restricts the attention of dynamics prediction to a dynamically-selected local neighborhood. We introduce a learning-based 3D region-of-interest (RoI) proposer that identifies a small subset of particles from the large graph that are likely to move. Assuming particles that are outside the RoI stay static during robot-terrain interaction improves speed by orders of magnitude in addition to higher dynamics prediction accuracy. L-GBND simultaneously learns the GBND dynamics and the RoI from simulation data, and both are fine-tuned on real data. Moreover, we introduce a novel boundary feature representation to address the problem that naïvely applying GBND to the identified subgraph leads to erroneous predictions. In our experiments, we verify both the efficiency and effectiveness of L-GBND for manipulating terrains of different materials.
II RELATED WORK AND DISCUSSION
@@
Model-Free Policy Learning. While model-free methods such as reinforcement learning (RL) and behavior cloning (BC) have achieved notable success in robotic manipulation [31, 32, 33, 34], they typically suffer from limited generalization, particularly in environments involving complex, high-dimensional, and long-horizon interactions like terrain deformation. These approaches require extensive task-specific data collection and tuning, often making them impractical for settings involving diverse or previously unseen terrain geometries. In the context of deformable terrain manipulation, the inability of model-free policies to extrapolate beyond the training distribution presents a fundamental limitation [35]. While such methods can learn reactive strategies within the training distribution, they lack an explicit mechanism for generalizing to new goals or terrain states. In contrast, dynamics model-based approaches retain predictive structure: given a learned forward model, planning can be performed zero-shot for unseen terrain states by simulating the effects of actions. This capability enables flexible and task-agnostic generalization, which is essential for terrain-aware manipulation. Because of these limitations, we do not consider model-free policies as baselines in this work. Our focus is exclusively on dynamics model learning, as it enables both general-purpose planning and scalability across diverse terrain interaction scenarios.
III Localized Graph-Based Neural Dynamics
@@
We present a framework for efficient and accurate terrain dynamics prediction under robot interaction, based on localized GBND (L-GBND) models. Our goal is to learn a computationally-efficient model to predict the shape of a terrain after robot-terrain interaction, such that L-GBND can be used in terrain manipulation planning. Specifically, letting $x_{t}$ denote the state of the terrain, we model $\hat{x}_{t+1}=f(x_{t},u_{t})$ where $u_{t}$ is the robot control and $t$ is the time index. We learn fine-grained prediction where the time step is 0.1 sec and $u_{t}$ indicates a small robot end-effector displacement. The dynamics model is trained on data recorded during three types of actions (scooping, pushing, and dumping), and we use a unified dynamics model for all actions. Pretraining is performed from simulation data and then the model is fine-tuned on real data.
@@
We choose L-GBND as our terrain model, which is a particle-based scene representation that provides flexibility in inter-particle interactions, an inductive bias toward locality in interactions, mass conservation, and preserves fine details of deforming solids and fluids [9, 11, 10]. To manage the high-dimensional state space of the terrain, we assume that only a small region of interest (RoI) $R_{t}(u_{t})\subset\mathbb{R}^{3}$ of the volume is affected by the interaction. Letting $p_{t}^{(i)}$ denote the $i^{\text{th}}$ particle in the terrain at time $t$, its future state $\hat{p}_{t+1}^{(i)}=p_{t}^{(i)}$ for any point $p_{t}^{(i)}\notin R_{t}(u_{t})$. Our method simultaneously learns $f$ and $R_{t}$. The RoI proposer is formulated as a super-level set of a learned function $g(p,u_{t})$ such that $g(p,u_{t})\geq 0$ if and only if $p\in R_{t}(u_{t})$. Using the RoI, we only predict the motion of a subgraph of the particle graph. The proposed framework is shown in Fig. 2.
@@
Our key innovations include: (1) learning $f$ and $g$ to minimize L-GBND memory usage and computation time while retaining high accuracy; and (2) introducing novel node features that enable learning L-GBND over arbitrary volumetric domains, ensuring that accuracy is preserved at the boundary of $R_{t}$, and that translation invariance is maintained.
III-A Graph-Based Scene Representation
We adopt a particle-based representation of the environment, in which both terrain and robot end-effectors are modeled as sets of particles. This representation is particularly suitable for learning physical dynamics in deformable and granular media, as each particle implicitly represents a small volume of matter. Terrain particles $P_{t}=\{p_{t}^{(i)}\}_{i=1}^{N_{p}}$ are instantiated from the observed terrain surface map, volumetrically down to a sufficiently large depth. Tool particles $T_{t}=\{p_{t}^{(i)}\}_{i=N_{p}+1}^{N_{p}+N_{t}}$ are sampled on the robot’s end-effector. For each particle, we connect it to the $k$ closest particles that meet a distance threshold using a nearest-neighbor algorithm, resulting in a sparse graph.
Follow prior work [10, 9, 11], we can define nodes in the graph as
where $\mathsf{v}_{t-H:t-1}^{(i)}$ is particle $p^{(i)}$’s velocity over the last $H$ time steps, $u_{t}^{(i)}$ is the impulse on the tool particle directly caused by robot action (displacement of the end-effector, hence, naturally zeros for terrain particles), and $c^{(i)}$ is the particle class (terrain or tool) embedding. The edge features for an edge $e_{t}^{(k)}$ connecting particles $i$ and $j$ are
III-B Graph-Based Neural Dynamics
@@
The dynamics of the particle system are modeled using a Graph Neural Network (GNN), which represents the environment state as a directed graph $x_{t}\equiv G_{t}(V_{t},E_{t})$ at each time step, where $V_{t}\triangleq\{v_{t}^{(i)}\}_{i=1}^{|V_{t}|}$ is the set of nodes, and $E_{t}\triangleq\{e_{t}^{(k)}\}_{k=1}^{|E_{t}|}$ is the set of edges, with each edge $e_{t}^{(k)}$ denoting a potential direct interaction from $v_{t}^{(i)}$ to $v_{t}^{(j)}$. Indirect interactions between particles are modeled in this framework using a message passing algorithm.
@@
The dynamics are defined as follows. Each node and edge are first encoded into a latent representation,
$z_{0}[v_{t}^{(i)}]=f_{V}^{\text{enc}}(v_{t}^{(i)}),$ $z_{0}[e_{t}^{(k)}]=f_{E}^{\text{enc}}(e_{t}^{(k)}).$ (3)
@@
Algorithm 1 L-GNBD
[…]
@@
  Predict $\tilde{g}$ based on $M_{t-H:t},u_{t}$ $\triangleright$ $R_{t}$ in Sec. III-C
@@
     $\emph{indices}\leftarrow\emph{indices}\cup\{i\}$
@@
Algorithm 2 MPPI Planning with L-GBND
@@
III-C Adaptive Fine-Grained Region of Interest Identification
Because the computation time and memory usage of GBND are closely tied to the number of particles and edges in the graph, limiting the graph’s size helps ensure efficiency. We learn an implicit function $g(p,u_{t})$ that defines the region of interest (RoI) $R_{t}(u_{t})=\{p\in\mathbb{R}^{3}\mid g(p,u_{t})\geq 0\}$.
We first extract the planar tool position and yaw from $u_{t}$ and define a 2D reference frame $T_{\text{tool}}$. We then predict a tool-centered 2D height function $\tilde{g}(x,y)$ such that $g(p,u_{t})=p_{z}-\tilde{g}(T_{\text{tool}}^{-1}(p_{x},p_{y}))$. In other words, points whose $z$ exceeds $\tilde{g}(x,y)$ (in the tool-centered frame) are included in the RoI.
Our approach uses a terrain heightmap and instantiates $\tilde{g}$ as a Convolutional Neural Network (CNN) architecture. A heightmap with grid resolution 2.5 cm is centered about the tool and cropped to a 0.4 m $\times$ 0.4 m region. At each time step, particles in the bounding box are squeezed into a $\times32$ heightmap, with each pixel concatenated with a 16D latent representation of the current control input. Our CNN is a U-Net-like [36] architecture with three levels of pooling, nonlinearity, and upsampling steps with skip connections. The output is a $\times16$ depth map representation of $\tilde{g}$.
Training uses simulation data. Moving particles are identified using a velocity threshold, and connected component filtering is applied to remove noise. The target heightmap is defined by combining the upper heightmap of non-moving particles (inflated to produce a bowl shape) with the bottom-up heightmap of moving particles. An $L_{1}$ loss is applied to non-empty cells, while a hinge loss at 0.5 m is used for empty cells. The pseudocode of L-GBND with scale reduction is presented in Algo. 1, and a typical simulation rollout with adaptive RoI scale reduction is presented in Fig. 3.
GBND GBND w/ $z$ features L-GBND
In Distribution 1.85 1.07 1.09
Out of Distribution 1.85 2.85 1.09
TABLE I: Comparison of node feature encodings for boundaries. GBND fails to generalize, and adding absolute height ($z$) overfits to training terrain elevation. L-GBND with normal features performs consistently across distributions.
III-D Boundary-Aware Node Features
GBND is best suited for tasks with similar training / test distributions, but our setting introduces the complication of unboundedness and varying regions of interest. As an example, past works using GBND for fluid and pile manipulation have used a fixed domain of particles, such that the network could implicitly learn the boundary conditions of a floor and walls [13, 12, 37, 18]. In our case, “boundary conditions” are not strictly-enforced boundaries, but are rather just more material that is unaffected by the current interactions due to energy dissipation.
In effect, particles adjacent to the RoI boundary should “feel” resistance as though the boundary were fixed. With naive node feature encoding (Equ. 1) as used in prior work, these particles will simply pass through the boundary as if there was nothing there. Alternatively, we experimented with adding “wall particles”, a boundary set of nodes, directly connected to nodes in the RoI, whose latent features are frozen. Not only does this increase memory usage due to massive wall particles, but in our experiments, inner particles were able to penetrate gaps in the sampled wall particles which introduces unnatural artifacts.
To encode appropriate boundary behavior, our approach incorporates an estimated geometric normal of particles within RoI $R_{t}$ [38]. In the interior, the normal information is not likely to help much, but normals are likely to be strongly aligned near the boundary. This allows our network to encode the effect that nodes near the RoI boundary should “feel” resistance in the direction of the normal. This encoding also has the favorable effect of translational invariance, which is that for the overall dynamics model $f$,
Experiments in Table I present validation losses of learned GBND models for different sets of node features. As baselines we compare standard GBND encoding as well as augmenting with particle height ($z$ features). In the in distribution case, the training and validation sets used terrains whose particles were sampled in a fixed $z$ range of [0 cm, 5 cm]. In the out of distribution case, the validation particle sets were randomly shifted in the $z$ axis by randomly sampled offsets from -100 to 100 m. Encoding the $z$ value significantly improves loss in the in-distribution case, demonstrating that the network is implicitly learning a “floor effect”. This implicit relationship is detrimental in the out-of-distribution case. In contrast, our L-GBND with proposed normal encoding leads to better generalization than the baseline in both regimes.
III-E Sim-to-Real Transfer
To transfer the pretrained L-GBND to the real world, we allow the robot to autonomously collect its own interaction data through self-play. Real-world trajectories are monitored using an overhead RGB-D camera, from which scene point clouds are reconstructed. Robot proprioception is used to remove robot particles and isolate in-tool particles by re-sampling within the tool’s convex hull. Since real-world data lacks explicit particle correspondences across time, we follow standard practices in the GBND literature and fine-tune the model using Chamfer Distance [39] and Earth Mover’s Distance [40] between predicted and observed point clouds. This enables effective learning without requiring explicit tracking of individual particles across frames.
III-F Terrain Manipulation Planning with Localized GBND
We use our L-GBND in a planner for terrain manipulation tasks. Our planner selects among parameterized scooping, pushing, and dumping trajectories to minimize the difference between the predicted terrain shape and a target heightmap. For scooping and pushing, we use the penetrate-drag-scoop (PDS) trajectory (Fig. 4) used in excavation [1, 41, 7], which maintains a 0 roll angle. This yields a 9D motion parameter space. Dumping uses a 6D parameterization (dumping location, yaw angle, pre- and post-dump pitch angles). Scooping and pushing are only allowed if the scoop is empty. The planner is based on Model Predictive Path Integral (MPPI) control [42]. A batch of trajectories is sampled in each iteration. The corresponding control sequences are rolled out with the dynamics model, and the outcomes are scored based on proximity to the goal shape. The best trajectory is selected and executed, and the terrain state is updated. Pseudocode is shown in Algo. 2.